• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Bartlett/Neil G-W

    Log entry time 12:42:22 on June26,2010

    Entry number 328388

    keyword=T8 trigger Restored

    We restored the 1024 Hz clock to the T8 trigger in both the scalers and Trigger supervisor. Run 1134 is a test run with only the T8 trigger with ps8=4, we saw 250 Hz.


    FIGURE 1